# Interacting Crystalline Topological Insulators in 2D with Spin-Orbit Coupling

### v1

@@ Abstract
Topology is routinely used to understand the physics of electronic insulators. However, for strongly interacting electronic matter, such as Mott insulators, a comprehensive topological characterization is still lacking. When their ground state only contains short range entanglement and does not break symmetries spontaneously, they generically realize crystalline symmetry protected topological phases (cFSPTs), supporting gapless modes at the boundaries or at the lattice defects. Here, we provide an exhaustive classification of cFSPTs in two-dimensions (2D) with $\mathrm{U}(1)$ charge-conservation and spinful time-reversal symmetries, namely those generically present in spin-orbit coupled insulators, for any of the 17 wallpaper groups. It has been shown that the classification of cFSPTs can be understood from appropriate real-space decorations of lower-dimensional subspaces, and we expose how these relate to the Wyckoff positions of the lattice. We find that all nontrivial one-dimensional decorations require electronic interactions. Furthermore, we provide model Hamiltonians for various decorations, and discuss the signatures of cFSPTs. This classification paves the way to further explore topological interacting insulators, providing the backbone information in generic model systems and ultimately in experiments.
@@ I Introduction
A classification of non-interacting FSPT phases, such as band insulators and superconductors, has been achieved in the tenfold way [1, 2, 3], where ten symmetry classes are defined depending on whether or not the single-particle Hamiltonians are symmetric under charge-conjugation symmetry, time-reversal symmetry (TRS), or a combination of the two. Over the years, weakly interacting electronic materials showcasing some of the boundary modes predicted by the tenfold way have been discovered: the paradigmatic examples include 111 The integer quantum Hall (IQH) states [4], while being the first realizations of fermionic topological phases of matter, are not protected by any symmetries owing to the fact that their gapless chiral boundary modes cannot be gapped even when all the symmetries are broken. IQH states are examples of invertible topological phases that are not SPT phases. These states carry long-range entanglement but do not support any nontrivial anyonic excitations [5]. The low-energy properties of invertible topological phases (SPT or not) are described by the so-called invertible topological field theories [6]. the quantum spin Hall effect [7, 8, 9], and the observation of gapless surface states in three-dimensional topological insulators [10, 11].
@@
In the tenfold way, the gapless boundary modes are protected by internal symmetries alone. However, the underlying lattice structure of crystalline matter naturally provides more symmetries such as translations, rotations, spatial inversion, or mirror operations. These symmetries enrich the known classification of FSPT phases by providing protection to distinct gapless boundary modes, which become gappable if the symmetry is broken. The corresponding phases are called crystalline FSPT (cFSPT) phases. Early examples of non-interacting cFSPT have been obtained by extending the tenfold way with crystalline symmetries [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], which led to the discovery of novel phases such as higher order topological phases [25, 26, 27, 28, 29]. A milestone in this direction were the development of the comprehensive frameworks of symmetry indicators [30, 31] and topological quantum chemistry [32, 33, 34, 35, 36], which allowed to characterize entire material databases. Since then, some of these phases have been observed in insulating electronic materials [37, 38, 39, 40, 41] and insulating or superconducting metamaterials [42, 43, 44].
@@
In the last decade, it has been realized that interaction effects can drastically change the classification of FSPTs, as well as cFSPTs. Following the seminal works of Fidkowski and Kitaev [45, 46], several works have shown that the classification of FSPTs (and cFSPTs) can be reduced due to the presence of interactions [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. In other words, nontrivial cFSPT phases may be connected to the trivial one by interaction terms. Parallel to these efforts, bosonic and fermionic SPT phases have been classified using the tools of algebraic topology such as cohomology and cobordism groups [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 6, 68]. In addition to reducing the non-interacting classification, interactions may also enhance the FSPT classification leading to intrinsically interacting phases that do not admit non-interacting counterparts [65, 69, 70, 71, 72, 6, 73].
@@
Despite the widespread presence of interactions in real electronic materials, the experimental realization and observation of interacting cFSPT phases have so far remained elusive, while some of them have been realized in cold atom experiments [86, 87]. A particularly relevant case is that of generic interacting electronic insulating materials with spin-orbit coupling and TRS. These properties encompass a range of materials, such as nonmagnetic Mott insulators. In the language of the tenfold way, such electronic insulators belong to the class AII, for which the relevant symmetries are $\mathrm{U}(1)$ charge-conservation and spinful TRS.
The classification of crystalline bosonic symmetry protected topological (SPT) states, as well as cFSPTs without U(1) conservation, with U(1) but without TRS, or with a subset of all possible spatial symmetries has been reported [78, 79, 80, 81, 82, 88]. However, a complete classification of interacting cFSPT phases with the symmetries of class AII and all space groups is so-far missing. In this work, we take a step forward in filling this gap by classifying all cFSPT phases in two-dimensional (2D) space in class AII, enriched by any of the 17 wallpaper groups. Using the real-space construction scheme described in Refs. 78, 79, 80, 81, 82, for each wallpaper group we tabulate the total classification and its decomposition in terms of classification of lower-dimensional FSPTs. Incidentally, we realize that Wyckoff positions [89] of a given wallpaper group provide a natural guideline for possible decorations, therefore bridging a well-known notion in the context of crystalline lattices with the concept of ‘unit cell subdivision’ adopted in Refs. 78, 79, 80, 81, 82. Interestingly, we find that all appropriate and nontrivial one-dimensional (1D) decorations at mirror symmetric lines correspond to intrinsically interacting cFSPT phases protected by the mirror symmetry. We provide toy-model Hamiltonians for both zero-dimensional (0D) and 1D decorations.
The rest of the paper is organized as follows. In Sec. II, we review the classification scheme we adopte, and we explicitly study the case of the honeycomb lattice, i.e., wallpaper group No. 17 (p6mm[89]. Therein, the classification for all wallpaper groups –our main result– is summarized in Table 1. Section III presents toy model Hamiltonians that can be used to decorate the points and lines of the 2D lattice. Section IV discusses various signatures of cFSPT phases. Finally in Sec. V we draw the conclusion and provide an outlook. Appendices A and B provide details on the classification of lower-dimensional decorations.
@@ II Classification by real-space construction
Our main result is summarized by Tab. 1, which lists the cFSPT classification in class AII for each of the 17 wallpaper groups. Many of the phases listed in the classification are intrinsically interacting (cyan entries), meaning that they can not be realized in the absence of interactions.
@@ 13 $p3$$\mathbb{Z}\times(\mathbb{Z}^{\,}_{3})^{3}\times{\color[rgb]{1,0,1}{\mathbb{Z}^{\,}_{2}}}$
Table 1: Classification of class AII cFSPTs for each of the 17 wallpaper groups. Each wallpaper group classification contains a $\mathbb{Z}$ factor corresponding to the total number of Kramers’ pairs in the unit cell, and a $\mathbb{Z}^{\,}_{2}$ factor for the intrinsic 2D SPT phase (the 2D “topological insulator”). The latter is colored in magenta to differentiate this phase which does not require crystalline symmetries nor interactions to be realized. The interaction enabled entries are highlighted in cyan, while the remaining entries are left in black.
@@ II.1 Overview of real-space construction
For a given wallpaper group, we construct cFSPTs by following the construction based on ‘decorating’ subspaces of the 2D space, which has been applied for classifying bosonic SPTs protected by point group symmetry, and cFSPTs in class A and class D of the tenfold way enriched by various crystalline symmetries [78, 79, 90, 80, 81, 82, 88]. In the following, we summarize the steps for the cFSPT construction through decoration, adapted to our case of class AII.
Unit cell subdivision.
First, we divide the 2D space into unit cells, which is allowed by the translation symmetries of the underlying lattice structure. We then subdivide each unit cell into $d$-dimensional subspaces, where $d=0,1,2$, which are invariant under some –or none, for the case of $d=2$– of the crystalline symmetries. Our construction differs from the previous work in Refs. 78, 79, 90, 80, 81, 82 by how we subdivide the unit cells. We obtain the relevant subspaces by considering the coordinates of the Wyckoff positions of the lattice (see Sec. II.3, for a review of Wyckoff positions and their relevant properties). If a given Wyckoff position has $d$ free parameters in its coordinate definition, we consider $d$-dimensional finite subspaces in the unit cell spanned by varying the free parameters of the coordinates. We refer to such subspaces as $d$-dimensional Wyckoff patches, where $d$ is equal to the number of free parameters in the definition of the Wyckoff position’s coordinates. This defines our $d$-dimensional subspaces, to be decorated by FSPT phases.
@@
A given $d$-dimensional Wyckoff patch may be left invariant under some of the crystalline symmetries such as rotations or reflections, which act internally within the Wyckoff patch. For each such $d$-dimensional Wyckoff patch, we enumerate all the possible $d^{\prime}$-dimensional FSPTs ($d^{\prime}\leq d$) that can be ‘glued’ to that patch, which we call decorations. The possible decorations are then $d^{\prime}$-dimensional FSPT phases protected by the internal symmetries of class AII, together with the crystalline symmetries that leave the corresponding Wyckoff patch invariant. We enumerate distinct decorations by assigning a topological index to each decoration. These indicate how the protecting symmetries are represented at a given Wyckoff patch. In Sec. II.2, we detail the classification of possible $d^{\prime}$-dimensional FSPT decorations. We remark that decorating a Wyckoff position with an FSPT does not imply that a physical degree of freedom (atom) needs to be placed there [see e.g., Fig. 3(c)].
@@ II.2 Classification of possible decorations
which comprises two subgroups. First, the subgroup $\mathrm{U}(1)^{\mathrm{F}}$ corresponds to the charge-conservation symmetry. The superscript $\mathrm{F}$ is to indicate that this subgroup contains the fermion parity symmetry $\mathbb{Z}^{\mathrm{F}}_{2}$ 222 If we parameterize the elements of $\mathrm{U}(1)^{\mathrm{F}}$ by $e^{\mathrm{i}\theta}$ with $\theta\in[0,2\pi)$, then fermion parity symmetry is generated by the element $\theta=\pi$. Fermion parity is a symmetry of any Hamiltonian built out of fermions and cannot be broken explicitly or spontaneously. . Second, the subgroup $\mathbb{Z}^{\mathrm{FT}}_{4}$ corresponds to the spinful TRS. Here too the superscript indicates that this subgroup contains fermion parity symmetry, since for spinful fermions TRS squares to fermion parity. To avoid double counting the fermion parity symmetry, we coset by the subgroup $\mathbb{Z}^{\mathrm{F}}_{2}$.
@@
The $\mathbb{Z}$ term in Eq. (3) is the classification of 0D FSPT phases protected only by charge-conservation and spinful TRS. Physically, it corresponds to the total number of Kramers’ pairs localized at the high-symmetry point, which we denote by the index $n\in\mathbb{Z}$. The remaining two terms $\mathbb{Z}^{\,}_{\mathrm{gcd}(2,a)}\times\mathbb{Z}^{\,}_{b}$ give the classification of 0D FSPT phases protected by the point group symmetries. Because of TRS, any 0D decoration can only carry real charges, i. e., ±$\pm 1$, under rotation and mirror symmetries. Whenever $\text{gcd}(2,a)=2$ or $b=2$, we define two $\mathbb{Z}^{\,}_{2}$-valued indices $r\in\mathbb{Z}^{\,}_{2}$ and $m\in\mathbb{Z}^{\,}_{2}$, which label the rotation and mirror charges, respectively, of a 0D decoration. Crucially, these $\mathbb{Z}^{\,}_{2}$ indices are absent in non-interacting electronic systems in AII, where wavefunctions can only transform trivially under the action of rotations or mirror transformations [91]. As we shall discuss in Sec. III, all 0D decorations with nontrivial indices $r$ or $m$ are intrinsically interacting.
@@
There are no crystalline symmetries in 2D space that leave a 2D decoration invariant. Thereofore, the total symmetry group for $d^{\prime}=2$ dimensional decorations is simply
@@ II.3 Decorating Wyckoff patches
Figure 1: (a) Wyckoff positions of the wallpaper group $p2mm$ (No. 6) with their labels and (c) illustration of the nontrivial site-symmetry groups of the lattice (either $C_{2v}$ or $C_{s}$ –marked in gray–). (b) Unit cell subdivision based on Wyckoff patches: 0D (circles), 1D (lines) and 2D (rectangles) patches are colored and labeled according to their corresponding Wyckoff position, following the color code of (a). (d) Example of 0D (dots), 1D (dashed lines) or 2D (rectangular regions) FSPTs decorations of the $p2mm$ lattice. Decorations are colored according to the Wyckoff position to which they belong, following the color code in (a). In all panels, a unit cell is marked by black lines.
We start this section by briefly recalling some useful concepts in the description of crystalline lattices. Every point in the lattice is characterized by its site-symmetry group, namely the finite subgroup of the space group transformations that leave the point invariant. A Wyckoff position is a set of points in the lattice whose site-symmetry groups are conjugate subgroups of the space group, i.e., the two site-symmetry groups are related by an element of the space group. Wyckoff positions are labeled by their multiplicity followed by a letter, and they are tabulated for all the space groups [89]. The ‘generic position’ on a 2D lattice is the Wyckoff position with trivial site-symmetry group The coordinates of the Wyckoff positions can either be fully determined or defined up to some free parameters.
@@
We now introduce the concept of Wyckoff patches: Given a Wyckoff position with $d$ free parameters in the definition of its coordinates, we call the $d$-dimensional region of the unit cell covered by varying the free parameters a $d$-dimensional Wyckoff patch.
By extension, we call the site-symmetry group of the Wyckoff patch the site-symmetry group of the Wyckoff position spanning the patch. Wyckoff positions with nontrivial site-symmetry groups lead to either 0D or 1D patches, while the general position results into a 2D patch. As Wyckoff positions provide a natural way to systematically characterize the lattice, we employ the concept of $d$-dimensional Wyckoff patches as our definition of unit cell subdivision.
As an example, Fig. 1(a) illustrates the Wyckoff positions of the unit cell of the wallpaper group No. 6 ($p2mm$) and Fig. 1(c) the corresponding site-symmetry group at each point. For instance, the explicit coordinates of the Wyckoff positions $1a$, $2e$ and $4i$ are
@@
where we expressed coordinates in terms of lattice vectors, namely $(x,y)=x\bm{a}_{x}+y\bm{a}_{y}$ with $\bm{a}_{x}=a\hat{\bm{x}},\,\bm{a}_{x}=b\hat{\bm{y}}$ ($a\neq b$), and the parameters $x,y$ are undetermined. From Eq. (7), we deduce that the $1a$, $2e$, and $4i$ Wyckoff positions result into 0D, 1D, and 2D Wyckoff patches, respectively. The $4i$ Wyckoff position is the general position of the lattice, it has trivial site-symmetry group and its coordinates span the full 2D space. From the Wyckoff patches, we identify the subdivision of the unit cell of $p2mm$, as shown in Fig. 1(b).
@@
With our newly introduced notion of unit cell subdivision, we can proceed and answer the question of how to decorate unit cells. A $d$-dimensional Wyckoff patch can be decorated by cFSPTs that: (i) are $d^{\prime}$-dimensional, with $d^{\prime}\leq d$, and (ii) are protected by the site-symmetry group of the Wyckoff patch and class AII.
@@
By appropriately combining the entries of Tab. 2, we can deduce the classification of the possible decorations on Wyckoff patches fulfilling (i) and (ii). The resulting classification is summarized in Tab. 3. The 0D Wyckoff patches can only be decorated by 0D cFSPTs, therefore the 0D entry of Tab. 2 exhausts their classification. The 1D Wyckoff patches can either be decorated by 1D cFSPTs, resulting into a $\mathbb{Z}^{\,}_{2}$ contribution to the classification, or by 0D cFSPTs, leading to an additional ×$\mathbb{Z}\times\mathbb{Z}^{\,}_{2}$ factor corresponding to the total number of Kramers’ pairs and total mirror eigenvalue of the 0D decorations, respectively. In fact, when multiple 0D decorations are stacked on a 1D line, they can be symmetrically and adiabatically moved along the line, and therefore the only relevant set of topological indices is the one counting their total charges per unit cell. The 2D Wyckoff patches do not have any crystalline symmetry acting internally, therefore the contribution of 2D decorations is given by the $\mathbb{Z}^{\,}_{2}$ graded 2D AII classification. Decorating these Wyckoff patches by 1D cFSPTs does not enrich the classification, as these must be trivial in the absence of additional internal symmetries, while 0D decorations contribute a $\mathbb{Z}$ index enumerating the total number of Kramers’ pairs. We label decorations at a Wyckoff patch by the topological indices introduced in Sec. II.2 with a subscript denoting the corresponding Wyckoff position. For instance, in the unit cell subdivision shown in Fig. 1, 0D FSPTs that can be glued to the 0D Wyckoff patch $1a$ are characterized by the triplet $n^{\,}_{1a}$, $r^{\,}_{1a}$, and $m^{\,}_{1a}$ that indicate the total number of Kramers’ pairs, their charge under rotations, and their charge under mirror symmetry, respectively. Figure 1(d) shows a possible set of decorations on the $p2mm$ lattice.
@@
Figure 2: (a) Wyckoff positions of the wallpaper group $p6mm$ (No. 17) and (b) mirror axes and rotation centers of the unit cell, corresponding to the nontrivial site-symmetry groups in the unit cell. (c)–(f) Representation of (c) 0D, (e)-(f) 1D and (d) 2D gappable decorations of Wyckoff patches, with the corresponding indices listed. In (e), we show an example of non-gappable decoration, and therefore with no index associated to it. (g),(h) Exemplification of the equivalence process $n^{\,}_{1a},n^{\,}_{2b}\rightarrow n^{\,}_{1a}-6k,n^{\,}_{2b}+3k$, described around Eq. (13) in the main text.

### v2

# Interacting Crystalline Topological Insulators in two-dimensions with Time-Reversal Symmetry

@@ Abstract
Topology is routinely used to understand the physics of electronic insulators. However, for strongly interacting electronic matter, such as Mott insulators, a comprehensive topological characterization is still lacking. When their ground state only contains short range entanglement and does not break symmetries spontaneously, they generically realize crystalline fermionic symmetry protected topological phases (cFSPTs), supporting gapless modes at the boundaries or at the lattice defects. Here, we provide an exhaustive classification of cFSPTs in two-dimensions with $\mathrm{U}(1)$ charge-conservation and spinful time-reversal symmetries, namely those generically present in spin-orbit coupled insulators, for any of the 17 wallpaper groups. It has been shown that the classification of cFSPTs can be understood from appropriate real-space decorations of lower-dimensional subspaces, and we expose how these relate to the Wyckoff positions of the lattice. We find that all nontrivial one-dimensional decorations require electronic interactions. Furthermore, we provide model Hamiltonians for various decorations, and discuss the signatures of cFSPTs. This classification paves the way to further explore topological interacting insulators, providing the backbone information in generic model systems and ultimately in experiments.
@@ I Introduction
A classification of noninteracting FSPT phases, such as band insulators and superconductors, has been achieved in the 10-fold way [1, 2, 3], where 10 symmetry classes are defined depending on whether or not the single-particle Hamiltonians are symmetric under charge-conjugation symmetry, time-reversal symmetry (TRS), or a combination of the two. Over the years, weakly interacting electronic materials showcasing some of the boundary modes predicted by the 10-fold way have been discovered: the paradigmatic examples include 111 The integer quantum Hall (IQH) states [4], while being the first realizations of fermionic topological phases of matter, are not protected by any symmetries owing to the fact that their gapless chiral boundary modes cannot be gapped even when all the symmetries are broken. IQH states are examples of invertible topological phases that are not SPT phases. These states carry long-range entanglement but do not support any nontrivial anyonic excitations [5]. The low-energy properties of invertible topological phases (SPT or not) are described by the so-called invertible topological field theories [6]. the quantum spin Hall effect [7, 8, 9], and the observation of gapless surface states in three-dimensional topological insulators [10, 11].
@@
In the 10-fold way, the gapless boundary modes are protected by internal symmetries alone. However, the underlying lattice structure of crystalline matter naturally provides more symmetries such as translations, rotations, spatial inversion, or mirror operations. These symmetries enrich the known classification of FSPT phases by providing protection to distinct gapless boundary modes, which become gappable if the symmetry is broken. The corresponding phases are called crystalline FSPT (cFSPT) phases. Early examples of noninteracting cFSPT have been obtained by extending the 10-fold way with crystalline symmetries [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], which led to the discovery of novel phases such as higher order topological phases [25, 26, 27, 28, 29]. A milestone in this direction was the development of the comprehensive frameworks of symmetry indicators [30, 31] and topological quantum chemistry [32, 33, 34, 35, 36], which allowed to characterize entire material databases. Since then, some of these phases have been observed in insulating electronic materials [37, 38, 39, 40, 41] and insulating or superconducting metamaterials [42, 43, 44].
@@
In the last decade, it has been realized that interaction effects can drastically change the classification of FSPTs, as well as cFSPTs. Following the seminal works of Fidkowski and Kitaev [45, 46], several works have shown that the classification of FSPTs (and cFSPTs) can be reduced due to the presence of interactions [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. In other words, nontrivial cFSPT phases may be connected to the trivial one by interaction terms. Parallel to these efforts, bosonic and fermionic SPT phases have been classified using the tools of algebraic topology such as cohomology and cobordism groups [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 6, 68]. In addition to reducing the noninteracting classification, interactions may also enhance the FSPT classification leading to intrinsically interacting phases that do not admit noninteracting counterparts [65, 69, 70, 71, 72, 6, 73].
@@
Despite the widespread presence of interactions in real electronic materials, the experimental realization and observation of interacting cFSPT phases have so far remained elusive, while some of them have been realized in cold atom experiments [86, 87]. A particularly relevant case is that of generic interacting electronic insulating materials with spin-orbit coupling and TRS. These properties encompass a range of materials, such as nonmagnetic Mott insulators. In the language of the 10-fold way, such electronic insulators belong to the class AII, for which the relevant symmetries are $\mathrm{U}(1)$ charge-conservation and spinful TRS.
The classification of crystalline bosonic symmetry protected topological (SPT) states, as well as cFSPTs without U(1) conservation, with U(1) but without TRS, or with a subset of all possible spatial symmetries has been reported [78, 79, 80, 81, 82, 88]. However, a complete classification of interacting cFSPT phases with the symmetries of class AII and all space groups is so-far missing. In this work, we take a step forward in filling this gap by classifying all cFSPT phases in two-dimensional (2D) space in class AII, enriched by any of the 17 wallpaper groups. Using the real-space construction scheme described in Refs. 78, 79, 80, 81, 82, for each wallpaper group we tabulate the total classification and its decomposition in terms of classification of lower-dimensional FSPTs. Incidentally, we realize that Wyckoff positions [89] of a given wallpaper group provide a natural guideline for possible decorations, therefore bridging a well-known notion in the context of crystalline lattices with the concept of ‘unit-cell subdivision’ adopted in Refs. 78, 79, 80, 81, 82. Interestingly, we find that all appropriate and nontrivial one-dimensional (1D) decorations at mirror symmetric lines correspond to intrinsically interacting cFSPT phases protected by the mirror symmetry. We provide toy-model Hamiltonians for both zero-dimensional (0D) and 1D decorations.
The rest of the paper is organized as follows. In Sec. II, we review the classification scheme we adopted, and we explicitly study the case of the honeycomb lattice, i.e., wallpaper group No. 17 (p6mm[89]. Therein, the classification for all wallpaper groups, our main result, is summarized in Table 1. Section III presents toy model Hamiltonians that can be used to decorate the points and lines of the 2D lattice. Section IV discusses various signatures of cFSPT phases. Finally in Sec. V we draw the conclusion and provide an outlook. Appendixes A and B provide details on the classification of lower-dimensional decorations.
@@ II Classification by real-space construction
Our main result is summarized by Table 1, which lists the cFSPT classification in class AII for each of the 17 wallpaper groups. Many of the phases listed in the classification are intrinsically interacting (cyan entries), meaning that they can not be realized in the absence of interactions.
@@ 13 $p3$$\mathbb{Z}\times(\mathbb{Z}^{\,}_{3})^{3}\times{\color[rgb]{1,0,1}{\mathbb{Z}^{\,}_{2}}}$
Table 1: Classification of class AII cFSPTs for each of the 17 wallpaper groups. Each wallpaper group classification contains a $\mathbb{Z}$ factor corresponding to the total number of Kramers’ pairs in the unit-cell, and a $\mathbb{Z}^{\,}_{2}$ factor for the intrinsic 2D SPT phase (the 2D “topological insulator”). The latter is colored in magenta to differentiate this phase which does not require crystalline symmetries nor interactions to be realized. The interaction enabled entries are highlighted in cyan, while the remaining entries are left in black.
@@ II.1 Overview of real-space construction
For a given wallpaper group, we construct cFSPTs by following the construction based on “decorating” subspaces of the 2D space, which has been applied for classifying bosonic SPTs protected by point group symmetry, and cFSPTs in classes A and D of the 10-fold way enriched by various crystalline symmetries [78, 79, 90, 80, 81, 82, 88]. In the following, we summarize the steps for the cFSPT construction through decoration, adapted to our case of class AII.
Unit-cell subdivision.
First, we divide the 2D space into unit-cells, which is allowed by the translation symmetries of the underlying lattice structure. We then subdivide each unit-cell into $d$-dimensional subspaces, where $d=0,1,2$, which are invariant under some, or none, for the case of $d=2$, of the crystalline symmetries. Our construction differs from the previous work in Refs. 78, 79, 90, 80, 81, 82 by how we subdivide the unit-cells. We obtain the relevant subspaces by considering the coordinates of the Wyckoff positions of the lattice (see Sec. II.3, for a review of Wyckoff positions and their relevant properties). If a given Wyckoff position has $d$ free parameters in its coordinate definition, we consider $d$-dimensional finite subspaces in the unit-cell spanned by varying the free parameters of the coordinates. We refer to such subspaces as $d$-dimensional Wyckoff patches, where $d$ is equal to the number of free parameters in the definition of the Wyckoff position’s coordinates. This defines our $d$-dimensional subspaces, to be decorated by FSPT phases.
@@
A given $d$-dimensional Wyckoff patch may be left invariant under some of the crystalline symmetries such as rotations or reflections, which act internally within the Wyckoff patch. For each such $d$-dimensional Wyckoff patch, we enumerate all the possible $d^{\prime}$-dimensional FSPTs ($d^{\prime}\leq d$) that can be “glued” to that patch, which we call decorations. The possible decorations are then $d^{\prime}$-dimensional FSPT phases protected by the internal symmetries of class AII, together with the crystalline symmetries that leave the corresponding Wyckoff patch invariant. We enumerate distinct decorations by assigning a topological index to each decoration. These indicate how the protecting symmetries are represented at a given Wyckoff patch. In Sec. II.2, we detail the classification of possible $d^{\prime}$-dimensional FSPT decorations. We remark that decorating a Wyckoff position with an FSPT does not imply that a physical degree of freedom (atom) needs to be placed there [see e.g., Fig. 3(c)].
@@ II.2 Classification of possible decorations
which comprises two subgroups. First, the subgroup $\mathrm{U}(1)^{\mathrm{F}}$ corresponds to the charge-conservation symmetry. The superscript $\mathrm{F}$ is to indicate that this subgroup contains the fermion parity symmetry $\mathbb{Z}^{\mathrm{F}}_{2}$222 If we parameterize the elements of $\mathrm{U}(1)^{\mathrm{F}}$ by $e^{\mathrm{i}\theta}$ with $\theta\in[0,2\pi)$, then fermion parity symmetry is generated by the element $\theta=\pi$. Fermion parity is a symmetry of any Hamiltonian built out of fermions and cannot be broken explicitly or spontaneously. Second, the subgroup $\mathbb{Z}^{\mathrm{FT}}_{4}$ corresponds to the spinful TRS. Here too the superscript indicates that this subgroup contains fermion parity symmetry, since for spinful fermions TRS squares to fermion parity. To avoid double counting the fermion parity symmetry, we coset by the subgroup $\mathbb{Z}^{\mathrm{F}}_{2}$.
@@
The $\mathbb{Z}$ term in Eq. (3) is the classification of 0D FSPT phases protected only by charge-conservation and spinful TRS. Physically, it corresponds to the total number of Kramers’ pairs localized at the high-symmetry point, which we denote by the index $n\in\mathbb{Z}$. The remaining two terms $\mathbb{Z}^{\,}_{\mathrm{gcd}(2,a)}\times\mathbb{Z}^{\,}_{b}$ give the classification of 0D FSPT phases protected by the point group symmetries. Because of TRS, any 0D decoration can only carry real charges, i. e., ±$\pm 1$, under rotation and mirror symmetries. Whenever $\text{gcd}(2,a)=2$ or $b=2$, we define two $\mathbb{Z}^{\,}_{2}$-valued indices $r\in\mathbb{Z}^{\,}_{2}$ and $m\in\mathbb{Z}^{\,}_{2}$, which label the rotation and mirror charges, respectively, of a 0D decoration. Crucially, these $\mathbb{Z}^{\,}_{2}$ indices are absent in noninteracting electronic systems in AII, where wave functions can only transform trivially under the action of rotations or mirror transformations [91]. As we shall discuss in Sec. III, all 0D decorations with nontrivial indices $r$ or $m$ are intrinsically interacting.
@@
There are no crystalline symmetries in 2D space that leave a 2D decoration invariant. Therefore, the total symmetry group for ($d^{\prime}=2$)-dimensional decorations is simply
@@ II.3 Decorating Wyckoff patches
Figure 1: (a) Wyckoff positions of the wallpaper group $p2mm$ (No. 6) with their labels and (c) illustration of the nontrivial site-symmetry groups of the lattice [either $C_{2v}$ or $C_{s}$ (marked in gray)]. (b) Unit cell subdivision based on Wyckoff patches: 0D (circles), 1D (lines), and 2D (rectangles) patches are colored and labeled according to their corresponding Wyckoff position, following the color code of (a). (d) Example of 0D (dots), 1D (dashed lines), or 2D (rectangular regions) FSPTs decorations of the $p2mm$ lattice. Decorations are colored according to the Wyckoff position to which they belong, following the color code in (a). In all panels, a unit-cell is marked by black lines.
We start this section by briefly recalling some useful concepts in the description of crystalline lattices. Every point in the lattice is characterized by its site-symmetry group, namely the finite subgroup of the space group transformations that leave the point invariant. A Wyckoff position is a set of points in the lattice whose site-symmetry groups are conjugate subgroups of the space group, i.e., the two site-symmetry groups are related by an element of the space group. Wyckoff positions are labeled by their multiplicity followed by a letter, and they are tabulated for all the space groups [89]. The “generic position” on a 2D lattice is the Wyckoff position with trivial site-symmetry group. The coordinates of the Wyckoff positions can either be fully determined or defined up to some free parameters.
@@
We now introduce the concept of Wyckoff patches: Given a Wyckoff position with $d$ free parameters in the definition of its coordinates, we call the $d$-dimensional region of the unit-cell covered by varying the free parameters a $d$-dimensional Wyckoff patch.
By extension, we call the site-symmetry group of the Wyckoff patch the site-symmetry group of the Wyckoff position spanning the patch. Wyckoff positions with nontrivial site-symmetry groups lead to either 0D or 1D patches, while the general position results into a 2D patch. As Wyckoff positions provide a natural way to systematically characterize the lattice, we employ the concept of $d$-dimensional Wyckoff patches as our definition of unit-cell subdivision.
As an example, Fig. 1(a) illustrates the Wyckoff positions of the unit-cell of the wallpaper group No. 6 ($p2mm$) and Fig. 1(c) the corresponding site-symmetry group at each point. For instance, the explicit coordinates of the Wyckoff positions $1a$, $2e$ and $4i$ are
@@
where we expressed coordinates in terms of lattice vectors, namely $(x,y)=x\bm{a}_{x}+y\bm{a}_{y}$ with $\bm{a}_{x}=a\hat{\bm{x}},\,\bm{a}_{x}=b\hat{\bm{y}}$ ($a\neq b$), and the parameters $x,y$ are undetermined. From Eq. (7), we deduce that the $1a$, $2e$, and $4i$ Wyckoff positions result into 0D, 1D, and 2D Wyckoff patches, respectively. The $4i$ Wyckoff position is the general position of the lattice, it has trivial site-symmetry group and its coordinates span the full 2D space. From the Wyckoff patches, we identify the subdivision of the unit-cell of $p2mm$, as shown in Fig. 1(b).
@@
With our newly introduced notion of unit-cell subdivision, we can proceed and answer the question of how to decorate unit-cells. A $d$-dimensional Wyckoff patch can be decorated by cFSPTs that: (i) are $d^{\prime}$-dimensional, with $d^{\prime}\leq d$, and (ii) are protected by the site-symmetry group of the Wyckoff patch and class AII.
@@
By appropriately combining the entries of Table 2, we can deduce the classification of the possible decorations on Wyckoff patches fulfilling (i) and (ii). The resulting classification is summarized in Table 3. The 0D Wyckoff patches can only be decorated by 0D cFSPTs, therefore the 0D entry of Tab. 2 exhausts their classification. The 1D Wyckoff patches can either be decorated by 1D cFSPTs, resulting into a $\mathbb{Z}^{\,}_{2}$ contribution to the classification, or by 0D cFSPTs, leading to an additional ×$\mathbb{Z}\times\mathbb{Z}^{\,}_{2}$ factor corresponding to the total number of Kramers’ pairs and total mirror eigenvalue of the 0D decorations, respectively. In fact, when multiple 0D decorations are stacked on a 1D line, they can be symmetrically and adiabatically moved along the line, and therefore the only relevant set of topological indices is the one counting their total charges per unit-cell. The 2D Wyckoff patches do not have any crystalline symmetry acting internally, therefore the contribution of 2D decorations is given by the $\mathbb{Z}^{\,}_{2}$ graded 2D AII classification. Decorating these Wyckoff patches by 1D cFSPTs does not enrich the classification, as these must be trivial in the absence of additional internal symmetries, while 0D decorations contribute a $\mathbb{Z}$ index enumerating the total number of Kramers’ pairs. We label decorations at a Wyckoff patch by the topological indices introduced in Sec. II.2 with a subscript denoting the corresponding Wyckoff position. For instance, in the unit-cell subdivision shown in Fig. 1, 0D FSPTs that can be glued to the 0D Wyckoff patch $1a$ are characterized by the triplet $n^{\,}_{1a}$, $r^{\,}_{1a}$, and $m^{\,}_{1a}$ that indicate the total number of Kramers’ pairs, their charge under rotations, and their charge under mirror symmetry, respectively. Figure 1(d) shows a possible set of decorations on the $p2mm$ lattice.
@@
Figure 2: (a) Wyckoff positions of the wallpaper group $p6mm$ (No. 17) and (b) mirror axes and rotation centers of the unit-cell, corresponding to the nontrivial site-symmetry groups in the unit-cell. (c)–(f) Representation of (c) 0D, (e), (f) 1D, and (d) 2D gappable decorations of Wyckoff patches, with the corresponding indices listed. In (e), we show an example of non-gappable decoration, and therefore with no index associated to it. (g), (h) Exemplification of the equivalence process $n^{\,}_{1a},n^{\,}_{2b}\rightarrow n^{\,}_{1a}-6k,n^{\,}_{2b}+3k$, described around Eq. (13) in the main text.